1 a) Write down, in words, the equation connecting speed, distance and time. [1] b) A car travels at a steady speed of 20 m/s. Calculate the distance travelled in 5
1
Expert's answer
2013-06-26T07:37:03-0400
1)
the equation connecting speed, distance and time:
V=s/t; s=V*t; t=s/V;
V-speed
t- time
s-distance
2)
A car travels at a steady speed of 20 m/s. Calculatethe distance travelled in 5 sec
s-?
Using formula mentioned above we obtains=V*t=20*5=100 m
[s]=[V/t]=[m*s/s]=[m]
Answer:
s=100 m